OpenStudy (anonymous):

Elimination : 2x+5y=-1 x+3y=0

OpenStudy (tkhunny):

What happens when you multiple the second equation by -2? -2x - 6y = 0 Why would I do that?

OpenStudy (anonymous):

oh it would be multiplyed by -2!

OpenStudy (tkhunny):

Now add the two equations.

OpenStudy (anonymous):

y=-1???

OpenStudy (anonymous):

then plug it in?

OpenStudy (anonymous):

No. When you add both equations you get: 2x+5y=-1 -2x-6y=0 ------------- -y=-1 so y=1 Then plug into either of your equations and solve for x.